# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 9.3. Invertibility of $\frac{1}{2}+\boldsymbol{K}$ on a cylinder
In this subsection, we keep the assumptions of the previous two subsections. In particular, we assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is on one side of its boundary, as usual (i.e., $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$, see Assumption 2.7). Let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$. To negate this statement, we shall write $\phi=0$ in AA.” Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
@@
Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}\big]\boldsymbol{h}^{\prime}=0$. Since $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity. Let us consider the indicial single layer potential $U:=\mathcal{S}_{0}\boldsymbol{h}^{\prime}$ (see Definition 9.2 with $\tau=0$). Proposition 9.5 then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}^{\prime}$ satisfy
@@
Let next $\Omega_{0}$ be a connected component of $\Omega_{-}^{\prime}$. Our assumptions give that $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$. Hence we obtain that $\boldsymbol{u}=0$ on $\Omega_{0}$, by Corollary 7.7. The same corollary gives that $p$ is constant in $\Omega_{0}$. Let us now show that this constant is actually 0. Indeed, the relation
@@
We have already observed that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{+}^{\prime}$. Proposition 9.6 then gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ of $\Omega^{\prime}$ (interior and exterior traces). Corollary 7.7 then gives that $\boldsymbol{u}=0$ on $\Omega^{\prime}$. Because $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on every connected component of $\Omega^{\prime}$, the same corollary gives that $p=0$ on $\Omega^{\prime}$. We have thus obtained that $U=0$ on both $\Omega_{\pm}^{\prime}$, and hence $U=0$ on $\mathfrak{S}$.
@@
The fact that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$ is also a classical result on pseudodifferential operators [16] (a proof can also be found in [63], Theorem 15.4.11, and in [67], Theorem 7.4). This completes the proof. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder $V$ and $V_{0}$ are smooth, translation invariant, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (6.1)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$, as usual. Let $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ be as in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$.
@@ 10.1. Invertibility of $\boldsymbol{S}$
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. Then $\widehat{\boldsymbol{S}\,}(0)$ is invertible and $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
@@
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, asymptotically translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The proof is the same as that of Proposition 10.1, but using Proposition 9.8 instead of Proposition 9.7. (As in that proof, the assumptions $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$ are needed for $\boldsymbol{\Xi}$ to be Fredholm, so that the layer potentials are defined.) ∎
@@
Let us assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$. We also assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Then $\boldsymbol{S}$ is an unbounded Fredholm operator on $L^{2}(\Gamma;TM)$. Equivalently, it is a (bounded) Fredholm operator $\boldsymbol{S}:H^{-1/2}(\Gamma;TM)\to H^{1/2}(\Gamma;TM)$.
@@
The operator ${\mathcal{R}_{\infty}}(\boldsymbol{S})$ is the single layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.9 are therefore satisfied by ${\mathcal{R}_{\infty}}(\boldsymbol{S})$, which is hence invertible. Because $\boldsymbol{S}$ is elliptic, Theorem 6.15 gives that $\boldsymbol{S}$ is Fredholm. ∎
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 (hence, in particular, $M$ is with straight cylindrical ends and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$). Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. We furthermore assume the following conditions.
@@
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
@@
To this end, let $\boldsymbol{h}\in H^{-1/2}(\Gamma;TM)$ be such that $\boldsymbol{S}\boldsymbol{h}=0$ and let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ be the corresponding single layer potential. The ellipticity of $\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and the elliptic regularity property of Theorem 8.16 assure that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Then, by Proposition 8.9, the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy the property
@@
Consequently, $\boldsymbol{S}$ is injective and, in view of the above arguments, it is invertible. Finally, the property that $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$ follows from Theorem 7.4 in [67] (see also Theorem 15.4.11 in [63]). ∎

### 28 May 2026 v2

@@ 9.3. Invertibility of $\frac{1}{2}+\boldsymbol{K}$ on a cylinder
In this subsection, we keep the assumptions of the previous two subsections. In particular, we assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and that $\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is on one side of its boundary, as usual (i.e., $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$, see Assumption 2.7). Let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$. To negate this statement, we shall write $\phi=0$ in AA.” Of course, in view of our conventions, the condition $\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
@@
Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}\big]\boldsymbol{h}^{\prime}=0$. Since $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by classical elliptic regularity. Let us consider the indicial single layer potential $U:=\mathcal{S}_{0}\boldsymbol{h}^{\prime}$ (see Definition 9.2 with $\tau=0$). Proposition 9.5 then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}^{\prime}$ satisfy
@@
Let next $\Omega_{0}$ be a connected component of $\Omega_{-}^{\prime}$. Our assumptions give that $\widehat{V}\not\equiv 0$ on $\Omega_{0}$. Hence we obtain that $\boldsymbol{u}=0$ on $\Omega_{0}$, by Corollary 7.7. The same corollary gives that $p$ is constant in $\Omega_{0}$. Let us now show that this constant is actually 0. Indeed, the relation
@@
We have already observed that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{+}^{\prime}$. Proposition 9.6 then gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ of $\Omega^{\prime}$ (interior and exterior traces). Corollary 7.7 then gives that $\boldsymbol{u}=0$ on $\Omega^{\prime}$. Because $\widehat{V}_{0}\not\equiv 0$ on every connected component of $\Omega^{\prime}$, the same corollary gives that $p=0$ on $\Omega^{\prime}$. We have thus obtained that $U=0$ on both $\Omega_{\pm}^{\prime}$, and hence $U=0$ on $\mathfrak{S}$.
@@
The fact that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$ is also a classical result on pseudodifferential operators [20] (see also Theorem 6.17; a proof can also be found in [83], Theorem 15.4.11, and in [87], Theorem 7.4). This completes the proof. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder $V$ and $V_{0}$ are smooth, translation invariant, and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (6.1)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$, see Theorem 6.17.
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$, as usual. Let $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ be as in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$.
@@ 10.1. Invertibility of $\boldsymbol{S}$
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\widehat{V}\succ 0$ on $M^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. Then $\widehat{\boldsymbol{S}\,}(0)$ is invertible and $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$.
@@
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, asymptotically translation invariant at infinity, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The proof is the same as that of Proposition 10.1, but using Proposition 9.8 instead of Proposition 9.7. (As in that proof, the assumptions $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$ are needed for $\boldsymbol{\Xi}$ to be Fredholm, so that the layer potentials are defined.) ∎
@@
Let us assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$. We also assume that $\widehat{V}\succ 0$ on $M^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Then $\boldsymbol{S}$ is an unbounded Fredholm operator on $L^{2}(\Gamma;TM)$. Equivalently, it is a (bounded) Fredholm operator $\boldsymbol{S}:H^{-1/2}(\Gamma;TM)\to H^{1/2}(\Gamma;TM)$.
@@
The operator ${\mathcal{R}_{\infty}}(\boldsymbol{S})$ is the single layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.9 are therefore satisfied by ${\mathcal{R}_{\infty}}(\boldsymbol{S})$, which is hence invertible. Because $\boldsymbol{S}$ is elliptic, Theorem 6.15 gives that $\boldsymbol{S}$ is Fredholm. ∎
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 (hence, in particular, $M$ is with straight cylindrical ends and $\widehat{V}\succ 0$ on $M^{\prime}$). Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. We furthermore assume the following conditions.
@@
    $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
@@
To this end, let $\boldsymbol{h}\in H^{-1/2}(\Gamma;TM)$ be such that $\boldsymbol{S}\boldsymbol{h}=0$ and let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ be the corresponding single layer potential. The ellipticity of $\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$, the ellipticity property of Theorem 8.16, and elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]) assure that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Then, by Proposition 8.9, the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy the property
@@
Consequently, $\boldsymbol{S}$ is injective and, in view of the above arguments, it is invertible. Finally, the property that $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$ follows from Theorem 6.17 (see Theorem 7.4 in [87] or Theorem 15.4.11 in [83]). ∎